Suppose that a consumer had a utility function given by: U=XY^3. This consumer has a budget of $12. Fill in the value of this consumer's demand function. X = Answer *(1/Px). (For instance, if X = 4/Px, then X = 4 * (1/Px) and enter 4)
Answer #1

U = XY3

Budget line: 12 = X.PX + Y.PY

Utility is maximized when MUX/MUY = PX/PY

MUX = \partial U/\partialX = Y3

MUY = \partial U/\partialY = 3XY2

MUX/MUY = (Y3) / (3XY2) = Y/3X = PX/PY

Y.PY = 3X.PX

Substituting in budget line,

12 = X.PX + 3X.PX

12 = 4X.PX

4X = 12 / PX

X = 3 / PX

X = 3 x (1/PX)

Required value = 3